Exercise 10.17. Claw. Jorge has three kids who spotted a Claw machine with toys they want. In stores the toys costs $10 each, but each play on the claw only costs $1. The probability of Jorge winning a game on the Claw is 0.12. Should he use the Claw to get the toys (he needs one toy per kid), or does he expect it to be cheaper to buy them in the stores?
Answer #1

expected number of trails to win a claw =1/p=1/0.12 =8.33

therefore expected value of toy if won by playing the claw =$ 8.33

he expect it to be cheaper if he use the Claw to get the toys